OpenStudy (dreamsawakes):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 4x + y2 − 6y = −4

OpenStudy (dreamsawakes):

@fuzzpuppy2 @Asad786

OpenStudy (dreamsawakes):

@zzr0ck3r

OpenStudy (dreamsawakes):

@ganeshie8

ganeshie8 (ganeshie8):

Start by grouping x terms and y terms separately

ganeshie8 (ganeshie8):

\(x^2 + 4x + y^2 − 6y = −4\) \((x^2 + 4x) + (y^2 − 6y) = −4\)

ganeshie8 (ganeshie8):

Do you know how to complete the square ?

OpenStudy (dreamsawakes):

._. I'm lost

OpenStudy (dreamsawakes):

When I see numbers, I completely zone out mostly @ equations

OpenStudy (dreamsawakes):

and them not being multiple choice hurts my soul

OpenStudy (hyuna301):

general ewq. of the circle is, \[x ^{2} + y ^{2} + 2gx + 2fy + c = 0\] centre = ( -g , -f) compare the above eq. with ur ques. and u will get the coordinates of the centre

ganeshie8 (ganeshie8):

can't avoid them, you should get used to it

OpenStudy (dreamsawakes):

x (x+4)+(y-6) y=-4

OpenStudy (dreamsawakes):

WAIT I THINK I GET IT

OpenStudy (dreamsawakes):

(x^2+4x+4)-4+(y^2-6y+9)-9=-4 (x+2)^2+(y-3)^2=9 (x+2)^2+(y-3)^2=3^2

OpenStudy (dreamsawakes):

?

OpenStudy (dreamsawakes):

the center of the circle is at (-2,3) and the radius is 3.

OpenStudy (dreamsawakes):

Am I right? @ganeshie8

OpenStudy (dreamsawakes):

@Hyuna301

OpenStudy (hyuna301):

yup.. ur are right

OpenStudy (dreamsawakes):

Thank you